# Approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search

### v1

@@ Abstract
We propose a variant of the approximate Bregman proximal gradient (ABPG) algorithm for minimizing the sum of a smooth nonconvex function and a nonsmooth convex function. Although ABPG is known to converge globally to a stationary point even when the smooth part of the objective function lacks globally Lipschitz continuous gradients, and its iterates can often be expressed in closed form, ABPG relies on an Armijo line search to guarantee global convergence. Such reliance can slow down performance in practice. To overcome this limitation, we propose the ABPG with a variable metric Armijo–Wolfe line search. Under the variable metric Armijo–Wolfe condition, we establish the global subsequential convergence of our algorithm. Moreover, assuming the Kurdyka–Łojasiewicz property, we also establish that our algorithm globally converges to a stationary point. Numerical experiments on $\ell_{p}$ regularized least squares problems and nonnegative linear inverse problems demonstrate that our algorithm outperforms existing algorithms.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is a continuously differentiable function, $g:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is a possibly non-differentiable convex function, and $\operatorname{cl}C$ is the closure of a nonempty open convex set $C\subset\mathbb{R}^{n}$. Optimization problems of the form (1.1) arise in various applications, including the maximum a posteriori (MAP) estimate [MAP, MAP2], ridge regression [ridge], the least absolute shrinkage and selection operator (LASSO) [LASSO]. In machine learning and signal processing, regularization or penalty terms are often introduced to prevent overfitting and impose the model structure. Some regularization terms are not necessarily differentiable.
@@
Numerous algorithms using proximal mappings have been proposed to solve (1.1). For instance, the proximal gradient method [Bruck1975-et, lions1979splitting, PASSTY1979383] and the fast iterative shrinkage-thresholding algorithm (FISTA) [fista] are included in the proximal algorithm. Convergence analysis of these algorithms with constant step-sizes typically rely on the global Lipschitz continuity of $\nabla f$, i.e., there exists $L>0$ such that $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|$ for any $x,y\in\mathbb{R}^{n}$. This condition is often restrictive and does not hold in certain applications in signal processing and machine learning.
@@
Bolte et al. [Bolte-2018] proposed the Bregman proximal gradient algorithm (BPG). This algorithm globally converges under the smooth adaptable property [Bolte-2018], also called relative smoothness [Lu-haihao], which is a relaxation of the global Lipschitz continuity of $\nabla f$. In recent years, the Bregman proximal gradient method has been improved from various perspectives. Hanzely et al. [Hanzely2021-sz] proposed accelerated Bregman proximal gradient algorithms for convex optimization problems using the triangle scaling property. Mukkamala et al. [Mukkamala2020-ea] proposed the accelerated version of BPG. Some researchers applied Bregman proximal-type algorithms to linear inverse problems [Bauschke2017-hg, Takahashi], nonnegative matrix factorization [Mukkamala2019-mk, Takahashi2026-rv], and blind deconvolution [Takahashi2023-uh].
@@
Since the subproblem of BPG is not always solved in closed form and is sometimes hard to solve depending on $\phi$, Takahashi and Takeda [Takahashi] proposed the approximate Bregman proximal gradient algorithm (ABPG), whose subproblem is easier to solve. Instead of the Bregman distance, ABPG uses the approximate Bregman distance (see also (3.2)), which is the second-order approximation of the Bregman distance. The subproblem of ABPG can be written by the sum of a quadratic formula and a regularizer. Moreover, if $\phi$ is separable, the subproblem of ABPG is reduced to $n$ independent one-dimensional optimization problems. ABPG uses the line search procedure to ensure the accuracy of the approximate Bregman distance. However, the global convergence of ABPG has not been established when $g\not\equiv 0$, and line search procedures lead to slow convergence in practice.
@@
where $x^{k}\in\operatorname{cl}C$, $\lambda>0$, and $\tilde{D}_{\phi}$ is the approximate Bregman distance (see, for its definition, (3.2)). The search direction of ABPG is defined by $d^{k}=y^{k}-x^{k}$, and ABPG searches $t_{k}\in(0,1]$ in each iteration to decide the step-size. The Armijo-like condition adopted in [Takahashi] imposes so stringent a condition on $t_{k}$ that it sometimes takes much time to calculate $t_{k}$ or makes $t_{k}$ very small to cause slow convergence. In this paper, we adopt a relaxed condition to ensure the validity for larger $t_{k}$. In addition to this, inspired by the Armijo–Wolfe-like condition, which Lewis and Overton [Lewis2013] applied to the quasi-Newton methods, we also propose the curvature condition for proximal algorithms. It aims to avoid excessively small step-sizes while ensuring that the search direction $d^{k}$ approaches 0 as $k\to\infty$. Similar to Bonettini et al. [Bonettini_2017], we also add a rule at the end of each iteration to select, as the updated point, the one that yields a smaller value of the objective function between $y^{k}$ and the point provided from the line search.
@@
The structure of this paper is as follows. Section 2 introduces essential notation such as the subdifferential, the Bregman distances, and the Kurdyka–Łojasiewicz property. In Section 3, we propose ABPG-VMAW and discuss its line search conditions. Section 4 shows properties of ABPG-VMAW and its global convergence. Section 5 presents numerical experiments on $\ell_{p}$ regularized least squares problems and nonnegative linear inverse problems. Finally, in Section 6, we present conclusions and future research directions.
@@
Let $B(\bar{x},r)=\{x\in\mathbb{R}^{n}\,|\,\|x-\bar{x}\|\leq r\}$ denote the ball of center $\bar{x}\in\mathbb{R}^{n}$ and radius $r>0$. Let $\operatorname{int}C$ and $\operatorname{cl}C$ be the interior and the closure of a set $C\subset\mathbb{R}^{n}$, respectively. The distance from a point $x\in\mathbb{R}^{n}$ to $C$ is defined by $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$. The indicator function $\delta_{C}$ is defined by $\delta_{C}(x)=0$ for $x\in C$ and $\delta_{C}(x)=+\infty$ otherwise. The sign function $\operatorname{sgn}(x)$ is defined by $\operatorname{sgn}(x)=-1$ for $x<0$, $\operatorname{sgn}(x)=0$ for $x=0$, and $\operatorname{sgn}(x)=1$ for $x>0$.
@@ 2.2 Bregman Distances
Definition 2.2 (Kernal Generating Distances [Bolte-2018, Definition 2.1]).
@@
For a kernel generating distancce $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\rightarrow\mathbb{R}_{+}$ is defined by
@@ 2.3 Kurdyka–Łojasiewicz Property
For $v>0$, we define $\Xi_{v}$ as a set of all continuous concave functions $\psi:[0,v)\rightarrow\mathbb{R}_{+}$ that is $\mathcal{C}^{1}$ on $(0,v)$ and satisfies $\psi(0)=0$, and whose derivative $\psi^{\prime}(x)$ is positive on $(0,v)$. We define the Kurdyka–Łojasiewicz property.
@@
Let $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be a proper and lower semicontinuous function. The function $f$ is said to satisfy the Kurdyka–Łojasiewicz property (for short: KL property) at $\bar{x}\in\operatorname{dom}\partial\Psi$ if there exist $v\in(0,+\infty]$, a neighborhood $U$ of $\bar{x}$, and a function $\psi\in\Xi_{v}$, such that for any $x\in U\cap[f(\bar{x})<f<f(\bar{x})+v]$, the following inequality holds:
@@
$\psi^{\prime}(f(x)-f(\bar{z}))\operatorname{dist}(0,\partial f(z))\geq 1.$ (2.6)
@@
    For any $x\in\operatorname{int}\operatorname{dom}\phi$ and $\lambda>0$, $\lambda g(u)+\frac{1}{2}\langle\nabla^{2}\phi(x)(u-x),u-x\rangle$ is supercoercive, that is,
@@
Assumption 3.1(i-iv) are standard assumptions for Bregman-type algorithms [Bolte-2018, Takahashi] and are generally satisfied in practice. For any $x\in C$, $\partial(g+\delta_{\operatorname{cl}{C}})(x)=\partial g(x)+\partial\delta_{\operatorname{cl}{C}}(x)=\partial g(x)$ holds because $x$ is an interior point of $C$ and $\operatorname{dom}g$ from Assumption 3.1(iii) and $\partial\delta_{\operatorname{cl}{C}}(x)=\{0\}$. For example, Assumption 3.1(v) holds if $\phi$ is strongly convex. Note that we will assume the strong convexity of $\phi$ in Assumption 4.1.
@@ 3.1 Approximate Bregman Proximal Gradient Algorithm
Let $\phi\in\mathcal{G}(C)$ be $\mathcal{C}^{2}$ on $C$. Takahashi and Takeda [Takahashi] define the approximate Bregman distance $\tilde{D}_{\phi}(u,x)\geq 0$, using the second-order approximation of $\phi(u)$ for $u\in\operatorname{dom}\phi$ around point $x\in\operatorname{int}\operatorname{dom}\phi$, as
@@
Assumption 3.2 ensures that the points generated by ABPG-VMAW are feasible. Obviously, when $C\equiv\mathbb{R}^{n}$, Assumption 3.2 holds. In the same discussion as [Bolte-2018, p. 2136] and [Takahashi, p.235], if $\phi$ is strongly convex, an envelope function $\inf_{u\in\operatorname{cl}{C}}\left\{\langle\nabla f(x),u-x\rangle+g(u)+\frac{1}{\lambda}\tilde{D}_{\phi}(u,x)\right\}$ is prox-bounded from [rockafellar2009variational, Exercise 1.24]. We have the following well-posedness result.
@@ 3.2 Variable Metric Armijo–Wolfe Line Search
Since $\Psi(y^{k})\leq\Psi(x^{k})$ is not necessarily guaranteed for the solution of the subproblem $y^{k}\in\tilde{\mathcal{T}}_{\lambda}(x^{k})$ with any $\lambda>0$, Takahashi and Takeda [Takahashi] introduced the line search procedure for ABPG. To ensure global convergence, we improved the condition of the line search procedure. The new condition is inspired by the line search method based on the Armijo–Wolfe condition proposed by Lewis and Overton [Lewis2013], and Miantao et al. [Miantao2024], and on the Armijo-like line search method introduced by Bonettini et al. [Bonettini_2017]. We also execute an update step to take a point corresponding to a smaller value of the objective function at the end of each iteration, inspired by Bonettini et al. [Bonettini_2017].
In order to define the search direction $d^{k}=y^{k}-x^{k}$, we solve the subproblem $y^{k}\in\tilde{\mathcal{T}}_{\lambda}(x^{k})$. Let $0<c_{1}<c_{2}<1$ and $\xi^{k}\in\partial g(x^{k})$. To ensure that $\Psi(x^{k+1})$ sufficiently decreases than $\Psi(x^{k})$, we impose the following condition on $t$:
@@
$\qquad<\Psi(x^{k})+c_{1}t\left(\langle\nabla f(x^{k}),d^{k}\rangle+g(x+d)-g(x)+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\right).$ (3.6)
To ensure that $x^{k+1}\in C$, we use $\delta_{\operatorname{cl}{C}}(x^{k}+td^{k})$. Takahashi and Takeda [Takahashi] used the sufficient decrease condition $\Psi(x^{k}+td^{k})<\Psi(x^{k})+c_{1}t(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k}))$, which is smaller than or equal to the right hand-side of (3.9) because of $\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle$. This fact implies that (3.6) allows a larger $t$ than existing conditions. Furthermore, to avoid excessively small step-sizes $t_{k}>0$, which could slow down convergence, we impose the condition given by
@@
$\quad-c_{1}t\left(\langle\nabla f(x^{k}),d^{k}\rangle+g(x+d)-g(x)+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\right),$ (3.9)
@@
Now we are ready to describe the proposed algorithm and its line search procedure for solving (1.1). The subproblem $\tilde{\mathcal{T}}_{\lambda}(x^{k})$ on line 1 is convex but strongly convex if $\phi$ is strongly convex (see also Assumption 4.1). To obtain the step-size $t_{k}$ satisfying $A_{k}(t_{k})<0$ and $W_{k}(t_{k})$ on line 1, we can use, for example, the bisection method (see, for more details, Section A). Moreover, $A_{k}(t_{k})<0$ implies $x^{k}+t_{k}d^{k}\in\operatorname{cl}{C}$ due to $\delta_{\operatorname{cl}{C}}(x^{k}+t_{k}d^{k})$ of $A_{k}(t_{k})$. Although we can choose any $\lambda>0$, it is better to use $\lambda<1/L$ for some $L$ as follows (see, for specific examples, Section 5).
@@
The parameter $\lambda$ can be any positive scalar. Note that the iteration number of the line search procedures would be large when $\lambda$ is large. In practice, it is better to choose $\lambda<1/L$, where $L>0$ is a parameter given by the smooth adaptable property, i.e., the pair $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad[Bolte-2018] if there exists $L>0$ such that both $L\phi-f$ and $L\phi+f$ are convex on $C$. The $L$-smad property provides the first-order approximation of $f$ by its descent lemma [Bolte-2018, Lemma 2.1]. Moreover, when $f$ and $\phi$ are $\mathcal{C}^{2}$, the pair $(f,\phi)$ is $L$-smad if and only if $L\nabla^{2}\phi(x)\preceq\nabla^{2}f(x)\preceq L\nabla^{2}\phi(x)$ holds for any $x\in C$. In order to achieve superior performance, it is recommended to choose a smaller $L$ and a $\phi$ that shares a similar structure to $f$. See, for more examples of the $L$-samd property, [Bolte-2018, Lemma 5.1][Bauschke2017-hg, Lemmas 7 and 8], [Takahashi, Proposition 24], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2026-rv, Theorem 4.1], [Takahashi2023-uh, Theorem 1], and [Dragomir2021-rv, Propositions 2.1 and 2.3].

### v2

@@ Abstract
We propose a variant of the approximate Bregman proximal gradient (ABPG) algorithm for minimizing the sum of a smooth nonconvex function and a nonsmooth convex function. ABPG is known to converge globally to a stationary point even when the smooth part of the objective function does not have a globally Lipschitz continuous gradient, and its iterates can often be expressed in closed form. However, ABPG relies on an Armijo line search to guarantee global convergence, which can slow down its practical performance. To address this issue, we propose a variant of ABPG with a variable metric Armijo–Wolfe line search. Under the variable metric Armijo–Wolfe condition, we establish global subsequential convergence of the algorithm. Moreover, assuming the Kurdyka–Łojasiewicz property, we also prove that the algorithm globally converges to a stationary point. Numerical experiments on $\ell_{p}$-regularized least squares problems and nonnegative linear inverse problems demonstrate that the proposed algorithm outperforms existing algorithms.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is a continuously differentiable function, $g:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is a possibly nondifferentiable convex function, and $\operatorname{cl}C$ is the closure of a nonempty open convex set $C\subset\mathbb{R}^{n}$. Optimization problems of the form (1.1) arise in various applications, including the maximum a posteriori (MAP) estimate [MAP, MAP2], ridge regression [ridge], the least absolute shrinkage and selection operator (LASSO) [LASSO]. In machine learning and signal processing, regularization or penalty terms are often introduced to prevent overfitting and impose the model structure. Some regularization terms are not necessarily differentiable.
@@
Numerous algorithms using proximal mappings have been proposed to solve (1.1). For instance, the proximal gradient method [Bruck1975-et, lions1979splitting, PASSTY1979383] and the fast iterative shrinkage-thresholding algorithm (FISTA) [fista] belong to the class of proximal algorithms. Convergence analysis of these algorithms with constant step-sizes typically relies on the global Lipschitz continuity of $\nabla f$, i.e., there exists $L>0$ such that $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|$ for any $x,y\in\mathbb{R}^{n}$. This condition is often restrictive and does not hold in certain applications in signal processing and machine learning.
@@
Bolte et al. [Bolte-2018] proposed the Bregman proximal gradient algorithm (BPG). This algorithm globally converges under the smooth adaptable property [Bolte-2018], also called relative smoothness [Lu-haihao], which is a relaxation of the global Lipschitz continuity of $\nabla f$. In recent years, the Bregman proximal gradient method has been improved from various perspectives. Hanzely et al. [Hanzely2021-sz] proposed accelerated Bregman proximal gradient algorithms for convex optimization problems using the triangle scaling property. Mukkamala et al. [Mukkamala2020-ea] proposed an accelerated version of BPG. Some researchers have applied Bregman proximal-type algorithms to linear inverse problems [Bauschke2017-hg, Takahashi], nonnegative matrix factorization [Mukkamala2019-mk, Takahashi2026-rv], and blind deconvolution [Takahashi2023-uh].
@@
Since the subproblem of BPG cannot always be solved in closed form and is sometimes hard to solve depending on the Bregman distance, Takahashi and Takeda [Takahashi] proposed the approximate Bregman proximal gradient algorithm (ABPG), whose subproblem is easier to solve. Instead of the Bregman distance, ABPG uses the approximate Bregman distance (see also (3.2)), which is the second-order approximation of the Bregman distance. The subproblem of ABPG can be written by the sum of a quadratic function and a regularizer. Moreover, if the Bregman distance is separable, the subproblem of ABPG is reduced to $n$ independent one-dimensional optimization problems. ABPG uses the line search procedure to ensure the accuracy of the approximate Bregman distance. However, the global convergence of ABPG has not been established when $g\not\equiv 0$, and line search procedures can lead to slow convergence in practice.
@@
where $x^{k}\in\operatorname{cl}C$, $\lambda>0$, and $\tilde{D}_{\phi}(u,x):=\frac{1}{2}\langle\nabla^{2}\phi(x)(u-x),u-x\rangle$ is the approximate Bregman distance with a twice continuously differentiable convex function $\phi$. The search direction of ABPG is defined by $d^{k}=y^{k}-x^{k}$, and ABPG searches for $t_{k}\in(0,1]$ in each iteration to decide the step-size. The Armijo-like condition adopted in [Takahashi] is so stringent that computing $t_{k}$ can be time-consuming, and it may force $t_{k}$ to be very small, which causes slow convergence. In this paper, we adopt a relaxed condition that allows larger values of $t_{k}$. In addition to this, inspired by the Armijo–Wolfe-like condition, which Lewis and Overton [Lewis2013] applied to the quasi-Newton methods, we also propose a curvature condition for proximal algorithms. It aims to avoid excessively small step-sizes while ensuring that the search direction $d^{k}$ approaches 0 as $k\to\infty$. Similar to Bonettini et al. [Bonettini_2017], we also add a rule at the end of each iteration to select, as the updated point, the one that yields a smaller value of the objective function between $y^{k}$ and the point provided from the line search.
@@
The structure of this paper is as follows. Section 2 introduces essential notation such as the subdifferential, the Bregman distances, and the Kurdyka–Łojasiewicz property. In Section 3, we propose ABPG-VMAW and discuss its line search conditions. Section 4 presents properties of ABPG-VMAW and its global convergence. Section 5 presents numerical experiments on $\ell_{p}$ regularized least squares problems and nonnegative linear inverse problems. Finally, in Section 6, we present conclusions and future research directions.
@@
Let $B(\bar{x},r)=\{x\in\mathbb{R}^{n}\,|\,\|x-\bar{x}\|\leq r\}$ denote the ball with center $\bar{x}\in\mathbb{R}^{n}$ and radius $r>0$. Let $\operatorname{int}C$ and $\operatorname{cl}C$ be the interior and the closure of a set $C\subset\mathbb{R}^{n}$, respectively. The distance from a point $x\in\mathbb{R}^{n}$ to $C$ is defined by $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$. The indicator function $\delta_{C}$ is defined by $\delta_{C}(x)=0$ for $x\in C$ and $\delta_{C}(x)=+\infty$ otherwise. The sign function $\operatorname{sgn}(x)$ is defined by $\operatorname{sgn}(x)=-1$ for $x<0$, $\operatorname{sgn}(x)=0$ for $x=0$, and $\operatorname{sgn}(x)=1$ for $x>0$.
@@ 2.2 Bregman Distances
Definition 2.2 (Kernel Generating Distances [Bolte-2018, Definition 2.1]).
@@
For a kernel generating distance $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\rightarrow\mathbb{R}_{+}$ is defined by
@@ 2.3 Kurdyka–Łojasiewicz Property
For $v>0$, we define $\Xi_{v}$ as a set of all continuous concave functions $\psi:[0,v)\rightarrow\mathbb{R}_{+}$ that are $\mathcal{C}^{1}$ on $(0,v)$ and satisfies $\psi(0)=0$, and whose derivative $\psi^{\prime}(x)$ is positive on $(0,v)$. We define the Kurdyka–Łojasiewicz property.
@@
Let $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be a proper and lower semicontinuous function. The function $f$ is said to satisfy the Kurdyka–Łojasiewicz property (for short: KL property) at $\bar{x}\in\operatorname{dom}\partial f$ if there exist $v\in(0,+\infty]$, a neighborhood $U$ of $\bar{x}$, and a function $\psi\in\Xi_{v}$, such that for any $x\in U\cap[f(\bar{x})<f<f(\bar{x})+v]$, the following inequality holds:
@@
$\psi^{\prime}(f(x)-f(\bar{x}))\operatorname{dist}(0,\partial f(z))\geq 1.$ (2.6)
@@
    For any $x\in\operatorname{int}\operatorname{dom}\phi$ and $\lambda>0$, $u\mapsto\lambda g(u)+\frac{1}{2}\langle\nabla^{2}\phi(x)(u-x),u-x\rangle$ is supercoercive, that is,
@@
Assumption 3.1(i)-(iv) are standard assumptions for Bregman-type algorithms [Bolte-2018, Takahashi] and are generally satisfied in practice. For any $x\in C$, $\partial(g+\delta_{\operatorname{cl}{C}})(x)=\partial g(x)+\partial\delta_{\operatorname{cl}{C}}(x)=\partial g(x)$ holds because $x$ is an interior point of $C$ and $\operatorname{dom}g$ from Assumption 3.1(iii) and $\partial\delta_{\operatorname{cl}{C}}(x)=\{0\}$. For example, Assumption 3.1(v) holds if $\phi$ is strongly convex. Note that we will assume the strong convexity of $\phi$ in Assumption 4.1.
@@ 3.1 Approximate Bregman Proximal Gradient Algorithm
Let $\phi\in\mathcal{G}(C)$ be $\mathcal{C}^{2}$ on $C$. Takahashi and Takeda [Takahashi] define the approximate Bregman distance $\tilde{D}_{\phi}(u,x)\geq 0$, using a second-order approximation of $\phi(u)$ for $u\in\operatorname{dom}\phi$ around point $x\in\operatorname{int}\operatorname{dom}\phi$, as
@@
Assumption 3.2 ensures that the points generated by ABPG-VMAW are feasible. Obviously, when $C\equiv\mathbb{R}^{n}$, Assumption 3.2 holds. In the same discussion as [Bolte-2018, p. 2136] and [Takahashi, p.235], if $\phi$ is strongly convex, the envelope function $\inf_{u\in\operatorname{cl}{C}}\left\{\langle\nabla f(x),u-x\rangle+g(u)+\frac{1}{\lambda}\tilde{D}_{\phi}(u,x)\right\}$ is prox-bounded from [rockafellar2009variational, Exercise 1.24]. We have the following well-posedness result.
@@ 3.2 Variable Metric Armijo–Wolfe Line Search
Since $\Psi(y^{k})\leq\Psi(x^{k})$ is not necessarily guaranteed for the solution of the subproblem $y^{k}\in\tilde{\mathcal{T}}_{\lambda}(x^{k})$ with any $\lambda>0$, Takahashi and Takeda [Takahashi] introduced the line search procedure for ABPG. To ensure global convergence, we improved the condition of the line search procedure. The new condition is inspired by the line search method based on the Armijo–Wolfe condition proposed by Lewis and Overton [Lewis2013] and Miantao et al. [Miantao2024], and on the Armijo-like line search method introduced by Bonettini et al. [Bonettini_2017]. We also execute an update step to take a point corresponding to a smaller value of the objective function at the end of each iteration, inspired by Bonettini et al. [Bonettini_2017].
In order to define the search direction $d^{k}=y^{k}-x^{k}$, we solve the subproblem $y^{k}\in\tilde{\mathcal{T}}_{\lambda}(x^{k})$. Let $0<c_{1}<c_{2}<1$ and $\xi^{k}\in\partial g(x^{k})$. To ensure that $\Psi(x^{k+1})$ is sufficiently smaller than $\Psi(x^{k})$, we impose the following condition on $t$:
@@
$\qquad<\Psi(x^{k})+c_{1}t\left(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k})+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\right).$ (3.6)
To ensure that $x^{k+1}\in C$, we use $\delta_{\operatorname{cl}{C}}(x^{k}+td^{k})$. Furthermore, to avoid excessively small step-sizes $t_{k}>0$, which could slow down convergence, we impose the condition given by
@@
$\quad-c_{1}t\left(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k})+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\right),$ (3.9)
@@
Now we are ready to describe the proposed algorithm and its line search procedure for solving (1.1). The subproblem $\tilde{\mathcal{T}}_{\lambda}(x^{k})$ on line 1 is convex, and it is strongly convex if $\phi$ is strongly convex (see also Assumption 4.1). To obtain the step-size $t_{k}$ satisfying $A_{k}(t_{k})<0$ and $W_{k}(t_{k})>0$ on line 1, we can use, for example, the bisection method (see, for more details, Section A). Moreover, $A_{k}(t_{k})<0$ implies $x^{k}+t_{k}d^{k}\in\operatorname{cl}{C}$ because of the term $\delta_{\operatorname{cl}{C}}(x^{k}+t_{k}d^{k})$ of $A_{k}(t_{k})$.
Remark 3.4 (Comparison with existing line search conditions).
For simplicity, we assume $C=\mathbb{R}^{n}$, so that $\delta_{\operatorname{cl}{C}}\equiv 0$. Our variable metric line search condition (3.6) allows a larger $t$ than the classical Armijo condition:
$\Psi(x^{k}+td^{k})$ (3.11)
$\qquad<\Psi(x^{k})+c_{1}t(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k}))$ (3.12)
$\qquad\leq\Psi(x^{k})+c_{1}t\left(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k})+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\right),$ (3.13)
where the second inequality follows from the variable metric $\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\geq 0$. When $\phi$ is strongly convex, the above second inequality holds with strict inequality for $d^{k}\neq 0$. Moreover, the parameter $\lambda$ can be chosen freely. As a practical guideline for choosing $\lambda$, we refer the reader to Remark 3.5.
On the other hand, our curvature condition (3.7), which uses $\xi^{k}$ on both sides, differs from the existing one. Although this may seem somewhat odd at first glance, it will play an important role in the proof of Lemma 4.9. When $g$ is nonsmooth, its subgradient typically fails to satisfy standard regularity conditions such as Lipschitz continuity, which is why $\xi^{k}$ appears on both sides.
Although we can choose any $\lambda>0$, it is better to use $\lambda<1/L$ for some $L$ as follows (see, for specific examples, Section 5).
@@
As already mentioned in Remark 3.4, the parameter $\lambda$ can be chosen to be any positive scalar. We now provide a practical guideline for selecting $\lambda$ to reduce the number of iterations. In practice, it is better to choose $\lambda<1/L$, where $L>0$ is a parameter given by the smooth adaptable property, i.e., the pair $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad[Bolte-2018] if there exists $L>0$ such that both $L\phi-f$ and $L\phi+f$ are convex on $C$. The $L$-smad property provides the first-order approximation of $f$ by its descent lemma [Bolte-2018, Lemma 2.1], which may reduce the number of iterations and computation time (see Figure 4). Moreover, when $f$ and $\phi$ are $\mathcal{C}^{2}$, the pair $(f,\phi)$ is $L$-smad if and only if $L\nabla^{2}\phi(x)\preceq\nabla^{2}f(x)\preceq L\nabla^{2}\phi(x)$ holds for any $x\in C$. In order to achieve superior performance, it is recommended to choose a smaller $L$ and a $\phi$ that shares a similar structure with $f$. See, for more examples of the $L$-smad property, [Bolte-2018, Lemma 5.1][Bauschke2017-hg, Lemmas 7 and 8], [Takahashi, Proposition 24], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2026-rv, Theorem 4.1], [Takahashi2023-uh, Theorem 1], and [Dragomir2021-rv, Propositions 2.1 and 2.3].